# Real-world Troublemaker: A Novel Track Testing Framework for Automated Driving Systems in Safety-critical Interaction Scenarios

### v1

@@ Abstract
Track testing plays a critical role in the safety evaluation of autonomous driving systems (ADS), as it provides real-world object targets and a safety-controllable interaction environment. However, existing track testing scenarios are often pre-fixed and limited, primarily due to the inflexibility of object target control methods and the lack of intelligent interactive behaviors. To overcome this limitation, we propose a novel track testing framework, Real-world Troublemaker, which can generate adversarial object target motion trajectories and facilitate intelligent interactions with the vehicle under test (VUT), creating a more realistic and dynamic testing environment. To enable flexible motion trajectories, cloud-controlled technology is utilized to remotely and dynamically control object targets to create a realistic traffic environment. To achieve intelligent interactions, an interactive concrete scenario generation method is introduced within a game-theoretic structure. In this approach, a novel quadratic risk utility function is designed to quantify maneuvering challenges, thereby promoting the diversity of interaction strategies. The dynamic interaction problem is reformulated as a Stackelberg game, allowing accurate reproduction of risk-based interactions. Furthermore, based on real risk interaction data, an important sampling-based strategy is proposed to increase the exposure frequency of safety-critical scenarios. The proposed framework has been successfully implemented at the Tongji University Intelligent Connected Vehicle Evaluation Base. Field test results demonstrate that Troublemaker can perform dynamic interactive testing of ADS accurately and effectively. Compared to traditional track testing methods, Troublemaker improves scenario reproduction accuracy by 65.2%, increases the diversity of target vehicle interaction strategies by approximately 9.2 times, and enhances exposure frequency of safety-critical scenarios by 3.5 times in unprotected left-turn scenarios.
@@
Automated driving systems, track testing, cloud-controlled object targets, safety-critical interaction scenarios.
@@ I Introduction
The safety of automated driving systems (ADS) must be ensured prior to their practical implementation, which requires a well-established testing framework. Existing test standards, such as ISO 26262 [1], UN R157 [2], and UN R171 [3], are not sufficient to comprehensively evaluate ADS. According to the driving automation levels defined by SAE J3016 from SAE International, a high-level ADS (i.e., Level 3 or higher) is expected to perform driving tasks independently and autonomously, with the driver no longer retaining continuous control over vehicle movement [4]. While ADS has already been deployed in various countries and regions, numerous automated vehicle traffic incidents highlight that safety testing for high-level ADS remains a critical technical challenge. In comparison to traditional vehicles and advanced driver assistance systems (ADAS), high-level ADS testing faces significant transformations and challenges, particularly in terms of both test subjects and requirements.
@@
In terms of test subjects, ADS has become more complex and intelligent, with the testing paradigm evolving from the traditional binary human-vehicle approach to a tightly coupled “vehicle-road-environment-task” system. From a testing entity perspective, ADS is no longer standalone like traditional vehicles but is now embedded within the broader traffic and transportation ecosystem [5]. With the rise of machine learning and end-to-end algorithms in high-level ADS technologies, the testing challenges have intensified. These advanced vehicles can outsmart predefined tests. As such, relying solely on dummy test equipment and modular testing methods is no longer sufficient to test these more intelligent and sophisticated systems.
@@
In terms of testing requirements, the primary focus of ADS has shifted from vehicle mechanical performance to driving performance in high-interaction scenarios. The ultimate goal of ADS testing is to validate the system’s performance under real-world road conditions, ensuring that ADS can outperform the average human driver in terms of safety [6]. The literature suggests that, to demonstrate that autonomous vehicles have a lower accident rate than human drivers in complex, real-world traffic environments, at least 17.7 billion kilometers of public road testing are required [7]. To fully capture the complexity and variability of these environments, the variables that define traffic scenarios are inherently complex, with interactive behaviors being random and diverse, which can lead to the curse of dimensionality [8]. Moreover, traditional testing methods may fail to address safety-critical scenarios, particularly due to the long-tail problem.
@@
Simulation testing, track testing, and public road testing are the three main pillars of ADS safety certification, each with its distinct advantages and limitations. Simulation testing effectively addresses the issue of mileage testing, but modeling precise and complex interaction scenarios remains a significant challenge [9]. Public road testing, while the most representative of real-world conditions, carries substantial and unacceptable risks, along with potential losses in the event of failures [10]. Additionally, public road testing is often inefficient, as many scenarios encountered are non-safety-critical and repetitive, which offers little value for the safety validation of ADS. Track testing presents a more controlled solution, providing a realistic yet manageable environment by utilizing soft targets. This enables the testing of interaction scenarios that closely simulate real traffic conditions [11]. In this setting, the vehicle under test (VUT) can be evaluated on roads with controlled interactions involving other actors, such as global vehicle targets and pedestrian targets, for crash testing. Significant progress has been made in both the equipment and the interaction scenario testing method.
@@
On the aspect of testing equipment, the definition of object targets for track testing systems has led to the establishment of corresponding standards. The ISO 19206 standard defines the test devices for targets and, through reasonable control methods, enables the actual ADS to be tested in traffic environments designed to simulate specific challenging driving conditions. However, the existing standards primarily focus on the physical characteristics of the test targets, such as size and reflective properties, while the control of these targets is still mainly oriented toward ADAS testing. On the aspect of interaction scenario testing methods, a widely accepted classification has been established, which includes the following:
@@ I-3 Search-based Methods
The state-of-the-art method in the field of ADS track testing is accelerated testing. Research focused on improving the efficiency of search-based testing methods. A risk scenario is defined as the combination of maneuver challenge and exposure frequency [22]. On the one hand, scholars have focused on improving the maneuver challenge. For instance, Koren et al. [23] proposed modeling complex interaction scenarios as a Markov decision process, incorporating randomness into the reward function and applying this model to a neural network to capture low-probability events. Althoff et al. [24] proposed a method to reduce the search space by altering the initial states of traffic participants to improve efficiency. Karimi et al. [25] introduced a framework for describing the complexity and modeled the generation of test cases as a constraint satisfaction search problem. On the other hand, Zhao et al. [26] proposed an accelerated lane-change scenario testing method based on the importance sampling theorem. The Feng team proposed a method combining importance sampling with key scenario-intensive reinforcement learning, theoretically proving that this approach can accelerate test efficiency by up to 2000 times under natural traffic flow on public roads [27] [28]. Accelerated testing methods can significantly shorten the testing mileage and time on closed-course tracks. However, these methods primarily adopt a hybrid approach where the AV operates in the real environment, while the object targets are generated through virtual simulations. This approach still faces challenges in modeling the exact and complex interaction scenarios.
@@
Based on the current state of research, it can be concluded that track testing for ADS faces two primary challenges, which also represent significant barriers. (i) The object targets used in track testing are limited by rigid control methods and low intelligence, often involving dummy targets against intelligent ADS. Additionally, the setup of these tests is costly, labor-intensive, and time-consuming. Existing test equipment mainly employs track-based or pre-set trajectory control methods, which are typically designed for ADAS testing and are not well-suited for high-level ADS testing [29]. (ii) Current testing methods are inefficient and lack interactivity. It is impractical to account for every potential scenario an automated vehicle could face within a fixed set of conditions, as the range of possible interactions is virtually limitless. However, most existing studies overlook the dynamic interactions between automated vehicles and other road users in complex traffic environments [30]. Furthermore, the majority of studies focus primarily on scenario design, with limited research on adversarial testing in real-world scenarios. Consequently, these testing methods fail to meet the growing need for ADS testing in interactive and dynamic environments.
@@
In response to the emerging demands and challenges in ADS testing, this paper proposes a novel testing framework, Troublemaker, designed for ADS testing on test tracks to address the issues identified earlier. The proposed framework leverages cloud-controlled technology to generate safety-critical test scenarios. Through real-time communication between object targets and the cloud control center, flexible control of multiple object targets is achieved, offering an efficient, automated, and cost-effective solution for ADS track testing. To fulfill the adversarial testing requirements for risk-interactive scenarios, a dynamic interaction game strategy is developed based on a quadratic risk utility function for maneuver challenge. Furthermore, importance sampling is introduced to enhance the exposure frequency of safety-critical scenarios, facilitating the accelerated testing of real interactive scenarios in a closed-course environment.
@@
  • Design a novel cloud-controlled test framework that enables dynamic interaction between object targets and the VUT, addressing limitations in flexibility, labor intensity, and setup time of traditional closed-course testing, while improving safety and efficiency.
  • Propose a dynamic game-based interaction model combined with an importance sampling strategy to enhance the exposure frequency to high-risk scenarios. This approach improves maneuver challenges and provides a solution to the inefficiency and uncontrollable risks in closed-course testing.
  • Demonstrate the “Troublemaker” testing system, the first cloud-coordinated object targets system for Level 3+ autonomous vehicles in closed-course testing. Deployed and validated in real-world tests at Tongji University, it facilitates large-scale automated testing and contributes to the advancement of closed-course testing for high-level autonomous vehicles.
@@ II Overview of the Troublemaker
As shown in Fig. 1, the overall architecture of Troublemaker is illustrated, which is a cloud-controlled object target testing system for test tracks. This framework consists of three main platforms: a cloud control platform, an object target motion platform, and a test-dedicated 5G network. Additionally, a test monitoring and command center, as well as a VUT test box (VBOX), have been developed for issuing test instructions and providing remote access to test personnel.
The cloud control platform is the core of the entire system and includes three primary modules: a data transmission protocol encoding/decoding module, a test scenario management module, and a test data storage module. The encoding/decoding module handles the data packaging process according to the defined VBOX/object target-cloud communication interface protocol and the cloud-monitoring center test instruction interface protocol. The test scenario management module is responsible for generating interactive test scenarios. It generates challenging and high-interaction scenarios for the VUT by planning the trajectories of object targets. The specific methodology is described in Section III. Additionally, real-time test data is uploaded to the cloud for storage and recording in the test data storage module. The communication network serves as the medium for system information exchange, using a dedicated 5G network to enable information interaction and data sharing. The object target platform, which are simulated traffic participants that interact with the VUT during testing, consists of a motion chassis and a vehicle shell. The motion chassis can execute trajectory commands issued by the cloud control platform and upload real-time motion state information to the cloud control platform. The test monitoring and command center, along with the VBOX vehicle collection device, are accessories designed to facilitate remote monitoring and issuing of test instructions. The VBOX is used to collect real-time state information of the VUT without the need for knowledge of the VUT’s algorithm or interface protocol.
@@
This paper focuses on the interaction concrete scenario generation method for object target control and interaction with the VUT, with an emphasis on the process of generating logical and concrete scenarios. According to the ISO 34501 standard, a logical scenario is described with the inclusion of parameters, whose values are defined as ranges. A concrete scenario is depicted with explicit parameter values, describing physical attributes. One way of generating a concrete scenario is to select a set of explicit parameters from the given logical scenario parameter space. In accordance with the standard’s definitions, this study will focus on the analysis of interaction scenario characteristics, the design of dynamic interaction strategies, and the enhancement of risk scenario exposure. The goal is to achieve the generation of interaction scenarios and their execution on-site.
Figure 1: Overview of the Troublemaker.
III-A Interaction Scenarios Character Analysis
Interaction risk scenarios occur when traffic participants simultaneously occupy road resources, leading to potential conflicts. In such scenarios, participants must consider both the constraints imposed by their counterpart’s decisions and the potential responses to their own decisions. Additionally, individual participants exhibit heterogeneous driving styles, and collision risks can take various forms. These scenarios are typical in real-world traffic, such as unprotected left turns at intersections and merging/diverging in roundabouts. The conditions for these scenarios are: (1) a potential conflict point exists between the participants, creating a risk scenario; (2) the interaction strategies of both parties increase the likelihood of a collision. Based on these features, a probabilistic model for the occurrence of specific interaction scenarios will be developed.
The generation of test scenarios aims to reproduce risk events within interaction scenarios, which can be used to efficiently test VUTs. According to the literature [9], an event of interest with VUTs is denoted as E (e.g., accident event). Our goal is to maximize the probability of conflict events occurring, which can be formulated as:
@@ III-A Interaction Scenarios Character Analysis
This specific scenario generation problem can be considered as designing a behavior policy that enables targets to aggressively interact with the VUT. Therefore, by selecting key interaction scenario parameters, a dynamic game approach can perform well in solving this issue. Intersections, as typical high-risk areas with frequent interactions in urban traffic environments, present even more challenges for automated driving systems, especially in unprotected left-turn scenarios [31]. Hence, we use this scenario as an example to introduce the testing method for interaction scenarios. For the purpose of this study, some assumptions are made.
@@
Remark 3: The Vehicle Under Test (VUT) exhibits generic driving behaviors, sharing common ”generic features,” such as maintaining safe distances and interacting safely with surrounding vehicles, as well as exhibiting driving regularity and smoothness.
Figure 2: Interactive critical scenario generation based on game strategy.
III-B Maneuver Challenge Improvement of Interactive Critical Scenarios
In this work, we develop a model to directly capture interactions based on a game-theoretic formulation, and design strategies for object target motion control, addressing the testing requirements from the perspective of enhanced interaction risks. Specifically, this includes the design of a risk interaction utility function, the development of a dynamic interaction strategy based on leader-follower game theory, and object target trajectory planning and control, as shown in Fig. 2.

### v2

# Real-world Troublemaker: A 5G Cloud-controlled Track Testing Framework for Automated Driving Systems in Safety-critical Interaction Scenarios

@@ Abstract
Track testing plays a critical role in the safety evaluation of autonomous driving systems (ADS), as it provides a real-world interaction environment. However, the inflexibility in motion control of object targets and the absence of intelligent interactive testing methods often result in pre-fixed and limited testing scenarios. To address these limitations, we propose a novel 5G cloud-controlled track testing framework, Real-world Troublemaker. This framework overcomes the rigidity of traditional pre-programmed control by leveraging 5G cloud-controlled object targets integrated with the Internet of Things (IoT) and vehicle teleoperation technologies. Unlike conventional testing methods that rely on pre-set conditions, we propose a dynamic game strategy based on a quadratic risk interaction utility function, facilitating intelligent interactions with the vehicle under test (VUT) and creating a more realistic and dynamic interaction environment. The proposed framework has been successfully implemented at the Tongji University Intelligent Connected Vehicle Evaluation Base. Field test results demonstrate that Troublemaker can perform dynamic interactive testing of ADS accurately and effectively. Compared to traditional methods, Troublemaker improves scenario reproduction accuracy by 65.2%, increases the diversity of interaction strategies by approximately 9.2 times, and enhances exposure frequency of safety-critical scenarios by 3.5 times in unprotected left-turn scenarios.
@@
Automated driving systems, track testing, 5G, cloud-controlled object targets, interaction scenarios.
@@ I Introduction
The safety of automated driving systems (ADS) must be ensured prior to their practical implementation, which requires a well-established testing framework [1]. Existing test standards, such as ISO 26262 [2], UN R157 [3], and UN R171 [4], are not sufficient to comprehensively evaluate ADS. According to the driving automation levels defined by SAE J3016 from SAE International, a high-level ADS (i.e., Level 3 or higher) is expected to perform driving tasks independently and autonomously, with the driver no longer retaining continuous control over vehicle movement [5]. While ADS has already been deployed in various countries and regions, numerous ADS traffic incidents highlight that safety testing for high-level ADS remains a critical technical challenge. In comparison to traditional vehicles and advanced driver assistance systems (ADAS), high-level ADS testing faces significant transformations and challenges, particularly in terms of both test subjects and requirements.
@@
In terms of test subjects, ADS has become more complex and intelligent, with the testing paradigm evolving from the traditional binary human-vehicle approach to a tightly coupled “vehicle-road-environment-task” system. From a testing entity perspective, ADS is no longer standalone like traditional vehicles but is now embedded within the broader transportation and traffic ecosystem [6]. With the rise of machine learning and end-to-end algorithms in high-level ADS technologies, the testing challenges have intensified. These advanced vehicles can outsmart predefined tests. As such, relying solely on dummy test equipment and modular testing methods is no longer sufficient to test these more intelligent and sophisticated systems.
@@
In terms of testing requirements, the primary focus of ADS has shifted from vehicle mechanical performance to driving performance in high-interaction scenarios. The ultimate goal of ADS testing is to validate the performance of the system under real-world conditions, ensuring that ADS can outperform the average human driver in terms of safety [7]. The literature suggests that, to demonstrate that autonomous vehicles have a lower accident rate than human drivers in complex, real-world traffic environments, at least 17.7 billion kilometers of public road testing are required [8]. To fully capture the complexity and variability of these environments, the variables that define traffic scenarios are inherently complex, with interactive behaviors being random and diverse, which can lead to the curse of dimensionality [9]. Moreover, traditional testing methods may fail to address safety-critical scenarios, particularly due to the long-tail problem.
@@
Currently, ADS assessment is performed mainly in simulation testing, track testing, and public road testing. Simulation testing effectively addresses the issue of mileage testing, but modeling precise and complex interaction scenarios remains a significant challenge [10]. Public road testing, while the most representative of real-world conditions, carries substantial and unacceptable risks, along with potential losses in the event of failures [11]. Additionally, public road testing is often inefficient, as many scenarios encountered are non-safety-critical and repetitive, which offers little value for the safety validation of ADS [12]. Compared with simulation and public road testing, track testing has its unique advantages. First, test tracks provide a realistic yet manageable environment by utilizing soft targets. Second, it enables the testing of interaction scenarios that closely simulate real traffic conditions [13]. In this setting, the vehicle under test (VUT) can be evaluated on roads with controlled interactions involving other actors, such as global vehicle targets and pedestrian targets, for crash testing. Third, dozens of test tracks are constructed around the world [14].
@@
Focused on ADS testing requirements, extensive research has been conducted around the test equipment and interaction scenarios test methods for track testing. On the aspect of test equipment, the definition of object targets for track testing systems has led to the establishment of corresponding standards. The ISO 19206 standard defines the test devices for targets and, through reasonable control methods, enables the actual ADS to be tested in traffic environments designed to simulate specific challenging driving conditions. However, the existing standards primarily focus on the physical characteristics of the test targets, such as size and reflective properties, while the control of these targets is still mainly oriented toward ADAS testing. On the aspect of interaction scenario test methods, a widely accepted classification has been established, which includes the following:
@@ I-3 Search-based Methods
The state-of-the-art method in the field of ADS track testing is accelerated testing. Research focused on improving the efficiency of search-based testing methods. A risk scenario is defined as the combination of maneuver challenge and exposure frequency [25]. On the aspect of improving the maneuver challenge. Koren et al. [26] proposed modeling complex interaction scenarios as a Markov decision process, incorporating randomness into the reward function and applying this model to a neural network to capture low-probability events. Karimi et al. [27] introduced a framework for describing the complexity and modeled the generation of test cases as a constraint satisfaction search problem. On the aspect of enhancing the exposure frequency, Zhao et al. [28] proposed an accelerated lane-change scenario testing method based on the importance sampling theorem. Feng et al. proposed a method combining importance sampling with key scenario-intensive reinforcement learning, theoretically proving that this approach can accelerate test efficiency by up to 2000 times under natural traffic flow on public roads [29] [30]. However, these methods primarily adopt a hybrid approach where the VUT operates in the test track, while other participants are generated through a virtual platform. This approach still faces challenges in modeling the exact and complex interaction scenarios.
@@
Although significant progress has been made in track testing, it faces two primary challenges, which also represent significant barriers. (i) The object targets used in track testing are limited by rigid control methods and low intelligence, often involving dummy targets against intelligent ADS. Additionally, the setup of these tests is costly, labor-intensive, and time-consuming. Existing test equipment mainly employs track-based or pre-set trajectory control methods, which are typically designed for ADAS testing and are not well-suited for high-level ADS testing [31]. (ii) Current testing methods are inefficient and lack interactivity. It is impractical to account for every potential scenario an automated vehicle could face within a fixed set of conditions, as the range of possible interactions is virtually limitless. However, most existing studies overlook the dynamic interactions between automated vehicles and other road users in complex traffic environments [32]. Furthermore, the majority of studies focus primarily on scenario design, with limited research on adversarial testing in real-world scenarios. Consequently, these testing methods fail to meet the growing need for ADS testing in interactive and dynamic environments.
@@
In response to the emerging demands and challenges in ADS testing, this paper proposes a novel testing framework, Real-world Troublemaker, designed for ADS testing on test tracks to address the issues identified earlier. The proposed framework leverages cloud-controlled technology to generate safety-critical test scenarios. Through real-time communication between object targets and the cloud control center, flexible control of multiple object targets is achieved, offering an efficient, automated, and cost-effective solution for ADS track testing. To fulfill the adversarial testing requirements for risk-interactive scenarios, a dynamic interaction game strategy is developed based on a quadratic risk utility function for maneuver challenge. Furthermore, importance sampling is introduced to enhance the exposure frequency of safety-critical scenarios, facilitating the accelerated testing of real interactive scenarios in a closed-course environment.
@@
  • Design a novel 5G cloud-controlled track test framework. The framework is capable of generating highly interactive scenarios, enabling remote control of the testing process. It addresses the limitations of existing test tracks, such as labor-intensive setups, time-consuming preparations, and poor dynamic interaction.
  • Propose an interactive testing scenario generation strategy, which designs a dynamic game model based on quadratic risk interaction utility to enhance maneuver challenges, along with a risk scenario exposure distribution function based on importance sampling. This provides a novel model to address the bottlenecks, such as low efficiency and difficulty in scenario reproduction.
  • Successfully deploy and validate a Troublemaker testing system, the first cloud-controlled track testing framework for ADS, contributing to the development of track testing for high-level ADS.
@@ II Overview of the Troublemaker
As shown in Fig. 1, the overall architecture of the Troublemaker is illustrated. This framework consists of three main platforms: a cloud control platform, an object target motion platform, and a test-dedicated 5G network. Additionally, a test monitoring and command center, as well as a VUT test box (VBOX), have been developed for issuing test instructions and providing remote access to test personnel.
The cloud control platform is the core of the entire framework and includes three primary modules: a data transmission protocol encoding/decoding module, a test scenario management module, and a cloud database. The encoding/decoding module handles the data packaging process according to the defined communication interface protocol and the cloud-monitoring center test instruction interface protocol. The test scenario management module is responsible for generating interactive test scenarios. It generates challenging and high-interaction scenarios for the VUT by planning the trajectories of object targets. The specific methodology is described in Section III. Additionally, real-time test data is uploaded to the cloud database for storage and recording. The communication network serves as the medium for system information exchange, using a dedicated 5G network to enable information interaction and data sharing. The object target platform, comprising mimic traffic participants that interact with the VUT, is composed of a motion chassis and a vehicle shell. The motion chassis is capable of executing trajectory commands from the cloud control platform while simultaneously uploading real-time motion state information to the platform. The test monitoring and command center, along with the VBOX vehicle collection device, are accessories designed to facilitate remote monitoring and issuing of test instructions. The VBOX is used to collect real-time state information of the VUT without the need for knowledge of the VUT’s algorithm or interface protocol.
Figure 1: Overview of the Troublemaker.
@@
In this section, we focus on the methods for generating interaction concrete scenarios. According to the ISO 34501 standard, a logical scenario is described with the inclusion of parameters, whose values are defined as ranges. A concrete scenario is depicted with explicit parameter values, describing physical attributes. In accordance with the definition of this standard, this study focuses on analyzing the characteristics of the interaction scenario, designing dynamic interaction strategies, and improving the exposure frequency of risk scenarios. The goal is to enable the generation of interaction concrete scenarios and their on-site execution.
III-A Interaction Scenarios Characteristics Analysis
Interaction risk scenarios occur when traffic participants simultaneously occupy road resources, leading to potential conflicts. In such scenarios, participants must consider both the constraints imposed by their counterpart’s decisions and the potential responses to their own decisions [33]. Additionally, individual participants exhibit heterogeneous driving styles, and collision risks can take various forms. These scenarios are typical in real-world traffic, such as unprotected left turns at intersections and merging/diverging in roundabouts. The conditions for these scenarios are: (1) a potential conflict point exists between the participants, creating a risk scenario; (2) the interaction strategies of both parties increase the likelihood of a collision. Based on these features, a probabilistic model for the occurrence of specific interaction scenarios will be developed.
The purpose of tracking testing is to reproduce risk events in interactive scenarios, which can be used to effectively test VUT. An event of interest with VUTs can be denoted as E (e.g., accident event). Our goal is to maximize the probability of conflict events occurring, which can be formulated as:
@@ III-A Interaction Scenarios Characteristics Analysis
This concrete scenario generation can be considered as designing a behavior policy that enables targets to aggressively interact with the VUT. Therefore, by selecting key interaction scenario parameters, a dynamic game approach can perform well in solving this issue. Intersections, as typical high-risk areas with frequent interactions in urban traffic environments, present even more challenges for automated driving systems, especially in unprotected left-turn scenarios [34]. Hence, we use this scenario as an example to introduce the testing method for interaction scenarios. For the purpose of this study, some assumptions are made.
@@
Remark 3: VUT exhibits generic driving behaviors, sharing common ”generic features,” such as maintaining safe distances and interacting safely with surrounding vehicles, as well as exhibiting driving regularity and smoothness.
Figure 2: Overview of interactive scenario generation based on dynamic game strategy.
III-B Maneuver Challenge Improvement Through a Dynamic Game Framework
We propose a game-theoretic model that directly captures interaction dynamics and designs object targets motion control strategies to address enhanced interaction risks. Specifically, this includes the design of a risk interaction utility function, the development of a dynamic interaction strategy based on leader-follower game theory, and object target trajectory planning and control, as shown in Fig. 2.